OpenStudy (anonymous):

What is the slope of the line? that passes through (1,5) and (-2,-4) a.-1/3 b.-3 c.1/3 d.3 @amistre64 @ParthKohli @myininaya

OpenStudy (anonymous):

@hartnn

OpenStudy (primeralph):



OpenStudy (anonymous):

would it be 1/3?

OpenStudy (neer2890):

slope of a line passing through two points (x1,y1) and (x2,y2) is \[m=\frac{ y2-y1 }{ x2-x1 }\]

OpenStudy (neer2890):

m=(y2-y1)/(x2-x1) so m=(5+4)/(1+2) m=9/3=3